OpenStudy (anonymous):

Simplify 25^((3)/(2)) I got 125 am I right?

Directrix (directrix):

Yes.

Directrix (directrix):

25^((3)/(2)) = [( 5^2) ] ^ (3/2) = [ 5 ^ (2 * (3/2)] = 5 ^ 3 = 125. Congrats!